OpenStudy (anonymous):

An automobile engine develops 133 hp(=99.18 kW) when rotating at 1820 rev/min. How much torque does it deliver?

OpenStudy (stormfire1):

In SI Units: \[kW=\frac{2 pi \tau (RPM)}{60,000}\]\[99.13kW=\frac{2 pi \tau(1820RPM)}{60,000}\]\[\tau=5201~N·m\] In British units: \[HP=\frac{2 pi \tau(RPM)}{33,000 ft-lb/min}\]\[133hp=\frac{2 pi \tau(1820RPM)}{33,000 ft-lb/min}\]\[\tau=383.8~lbf·ft\]

OpenStudy (stormfire1):

err...that should have been 520.1Nm The decimal disappeared I swear...

OpenStudy (anonymous):

Could you please show me the formula where torque is expressed in other terms?

OpenStudy (anonymous):

@stormfire1

OpenStudy (anonymous):

Oh, I sorted it out.

OpenStudy (stormfire1):

ok :)

OpenStudy (anonymous):

Thanks a lot, buddy! One more question, how to derive the equation you gave me? @stormfire1

OpenStudy (anonymous):

I mean this equation \[kW=\frac{2 \pi \tau (RPM)}{60,000}\]

OpenStudy (stormfire1):

It's essentially this equation:\[W=\tau 2piRPS\] converted to RPM See: http://en.wikipedia.org/wiki/Torque

OpenStudy (stormfire1):

There's a more complete derivation on that site

OpenStudy (anonymous):

Oh, next time I'll check wikipedia first :))) Nevertheless, thanks a lot!)))

OpenStudy (stormfire1):

gotta love wiki :P